A ball is dropped from height h . How much time taken by the ball to reach the ground (take g as acceleration due to gravity) ?

Grade:9

2 Answers

Tushar Kumar Verma
7 years ago
Initial velocity,u=0by using newton`s third equation of motion h=ut+(1/2gt^2),we get-h=1/2gt^22h=gt^2t=square root (2h/g)
PASUPULETI GURU MAHESH
7 years ago
intial U =0
we know that h=ut+1/2 at2
                     h=½ gt2
                      t=root of (2h/g)
time taken ball reach the ground is t=root of (2h/g)
   THANK YOU